# Inductive-Deductive Strategy Reuse for Multi-Turn Instructional Dialogues

### v1

@@ 6 Related Work
Instructional Dialogue Generation.
Aligning LLMs with human expectations requires massive high-quality instructions for fine-tuning LLMs. With the success of top-tier LLMs (e.g., GPT-4), it is feasible to use these strong LLMs to automatically construct a variety of instructions, e.g., single-turn instructions Wang et al. (2023b); Wu et al. (2023); Anand et al. (2023); Honovich et al. (2023); Yu et al. (2023); Xu et al. (2023a) and multi-turn instructions Ding et al. (2023); Xu et al. (2023b); Li et al. (2023); Ji et al. (2023). The latest approach is leveraging two role-played LLMs to interact: one simulating a user to pose instructions, and the other acting as a system agent to respond. These two LLMs are provided corresponding prompts to produce outputs. In contrast, our approach introduces a trainable user simulator that leverages instructional strategies to generate high-quality instructions.
Learn to Pose Instructions.
Kong et al. (2023a) and Sun et al. (2023) propose to train user simulators for posing instructions. Besides, some works in similar tasks study how to ask high-quality questions, including question generation in open-domain dialogue systems Wang et al. (2018); Shen et al. (2021) and conversational question generation Gu et al. (2021); Li et al. (2022); Zeng et al. (2023). These works model the dialogue flow in multi-turn dialogues without explicit guidance for posing high-quality instructions and questions. The difference is that IDEAS introduces instructional strategies to explicitly model the dialogue flow.
@@
Some researches Ou et al. (2022); Bao et al. (2023) utilizes a sequence of knowledge pieces (e.g., keywords, Wikipedia pages) to simulate the trajectory of dialogue flows for guiding open-domain dialogue generation. Compared to existing works, we are the first to leverage abstract knowledge, i.e., high-level instructional strategies, for dialogue flow modeling. We posit that LLMs have the capability to identify keywords and related information within extensive corpora. However, LLMs encounter difficulty in deriving high-level strategies. Therefore, introducing these abstract strategies could potentially serve as a more effective direction, supporting the model in the generation of superior multi-turn dialogues.
@@
Our work is based on inductive learning Michalski (1983), which has shown promising results in various scenarios, including reasoning tasks Wang et al. (2023a); Zhu et al. (2023); Sun et al. (2024) and autonomous agents Park et al. (2023); Zhao et al. (2023). However, our work focuses on inducing high-level instructional strategies from human-machine dialogues and then applying them to guide instruction generation in new dialogue scenarios, which has not been explored in instructional dialogues.
@@ 7 Conclusion
This paper presents an inductive-deductive strategy reuse method, IDEAS, to generate diverse, in-depth, and insightful instructions for building multi-turn instructional dialogues. Specifically, IDEAS first induces high-level instructional strategies from various human-machine dialogues via inductive reasoning. Afterward, IDEAS deductively applies these strategies to new dialogue scenarios, where the instructional strategies explicitly guide the dialogue flow and facilitate high-quality instructions. Experimental results show that IDEAS can generate diverse, in-depth, and insightful instructions. The constructed instructional dialogues can be further used to improve the performance of the downstream chat model. We hope IDEAS will inspire more human-inspired methods to build more high-quality multi-turn instructional dialogues.
@@
Due to limited computational and financial resources, we only select 13B pre-trained language models, i.e., LLaMA-2, for training the user simulator. Recent research suggests that LLMs exhibit emerging capabilities when they expand beyond a certain threshold Wei et al. (2022a). IDEAS implemented based on larger LLMs may bring different effects. However, we do not have enough resources to conduct relevant experiments. We welcome further researchers to study the benefits of abstract knowledge, i.e., instructional strategies, to larger LLMs.
@@
We design instructional strategy extraction and generalization, along with quality control by prompting GPT-4 to implement. We also choose GPT-4 as the system agent. However, GPT-4 is only accessed through a programming interface that essentially adds a black box on top of a black box. The mechanisms behind these interfaces may change at any time, so the constructed dialogues from different periods may change.
@@
Given that GPT-4 is trained on online data, it is conceivable that GPT-4 may encode pervasive biases that perpetuate stereotypes, discrimination, or marginalization of specific languages or communities. This results in potentially toxic and harmful answers from GPT-4 as the system agent. In addition, our evaluation process involves manual intervention by two professional annotators. Each annotator is compensated $0.2$ per instance.
Appendix A Task Definition
Instruction Generation.
This generative model is fine-tuned on real human-machine dialogues as a user simulator, which iteratively interacts with a system agent (e.g., GPT-4) to obtain multi-turn instructional dialogues for different dialogue scenarios.
Response Generation.
Appendix B IDEAS
B.1 Model Training
@@
The BERT-based ranker may have limited performance. Therefore, this may raise a question: why not train a stronger ranker to select the top-$1$ strategy or randomly select a strategy from the top-$k$ strategies, i.e., the candidate, to avoid further selection by the user simulator? This is mainly due to the following reasons: (1) There are generally multiple appropriate instructional strategies for a given dialogue history, so it is unreasonable to choose the top-$1$ strategy. (2) It is reasonable for the user simulator to sample one from multiple appropriate strategies in the candidate. (3) Limited performance means that not every strategy in the candidate is appropriate, so random selection from the candidate is also unreasonable. (4) The strategy selection of the user simulator is better than that of the ranker, whose secondary selection can select more appropriate strategies. (5) It is easy for the user simulator to select a strategy before generating instructions, but training a stronger ranker will greatly increase the complexity of IDEAS. Overall, this lightweight model is sufficient for a coarse selection of candidates.
Appendix C Experimental Details
C.1 Data
For instruction generation, we use the human-machine instructional dialogues from ShareGPT  [^2] https://huggingface.co/datasets/shareAI/ShareGPT-Chinese-English-90k  [^3] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna
@@ C.1 Data
to obtain instructional strategies, along with training the user simulator and the ranker in the induction stage. We first clean the dialogues using the provided method  [^4] https://huggingface.co/datasets/shareAI/ShareGPT-Chinese-English-90k/blob/main/shareGPT/filter_data.py and obtain $56,929$ dialogues. Subsequently, we process each multi-turn dialogue $\{\boldsymbol{q}_{0},\boldsymbol{a}_{0},\dots,\boldsymbol{q}_{T},\boldsymbol{a}_{T}\}$ into $T$ pairs of dialogue histories and instructions, e.g., <$<\{\boldsymbol{q}_{0},\boldsymbol{a}_{0},\dots,\boldsymbol{a}_{t-1}\},\boldsymbol{q}_{t}>$. We further obtain $211,057$ pairs and the corresponding original instructional strategies. Afterward, we achieve $1,593$ high-level instructional strategies via instructional strategy abstraction. Furthermore, we randomly sample $200$ pairs of dialogue histories and instructions and employ three human annotators to evaluate the appropriateness of original and high-level instructional strategies. About $95.67\%$ original instructional strategies are accepted by at least two annotators, and about $95\%$ high-level instructional strategies are accepted. Similarly, we evaluate the accuracy of the original strategy abstraction, and at least two annotators accept about $94.33\%$ corresponding high-level strategies.
In the deduction stage, we randomly sample $10,000$ dialogues from unused ShareGPT and UltraChat datasets respectively, and then intercept the first-turn dialogues $\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}\textquoteright_{0}\}$ as opening lines. We set the maximum interaction to $10$ rounds and the maximum of regeneration in quality control to $5$. The average number of regeneration is $0.69$. We also evaluate the proportion of instructional strategies selected by the user simulator from the candidates, which is $97.76\%$. Besides, we evaluate the appropriateness of the selected strategies by the user simulator, and the annotators accept about $93.82\%$ selected instructional strategies. Furthermore, we manually evaluate the correctness of quality control, and at least two annotators accepted $93.33\%$ of the judgment results. After achieving our constructed instructional dialogues $\mathcal{D}_{chat}$, we clean dialogues from $\mathcal{D}_{chat}$ by removing subsequent rounds starting from the round in which the error content appears. The error content contains the empty instructions and the incorrect answers caused by exceptions in calling the interface API of the system agent. Consequently, the average turn in the generated dialogues is $9.33$, and the average length of instructions and answers are $36.90$ and $654.04$ respectively.
C.2 Implementation Details
@@
We prompt gpt-4-1106-preview for instructional strategy extraction and abstraction, along with quality control. We set the temperature to 00, the presence_penalty to $0.6$, and the frequency_penalty to 00. For strategy clustering, the threshold $\epsilon$ is set to $0.5$. The entire cluster process is implemented by Faiss vectorstore Johnson et al. (2019). For ranking, we set the threshold $\eta$ to $0.5$. We limit the size of the candidates to 50 due to the context length limitation. When generating instructions, we use FasterTransformer  [^5] https://github.com/NVIDIA/FasterTransformer for inference acceleration. And the temperature is set to $0.7$, the top-p is $0.9$, and the max new tokens is $96$. The average inference time for one instruction is approximately $0.2$ second. Besides, we employ gpt-4-1106-preview as the system agent, where the parameters are default.
@@
Prompt for Instructional Strategy Extraction
You are very good at learning from real-human dialogues about how they generate high-quality instructions. The dialogue history is: {Dialogue History}. The next instruction generated by the human is: {Instruction}. Please analyze why humans generate such instruction based on the above dialogue history, and give the instructional strategy you have learned. Good instructional strategies need to meet the following requirements:
(1) Instructional strategies should be expressed in simple phrases;
(2) Instructional strategies should not include specific information from the dialogue history. To better assist you in inducing instructional strategies, I will provide you with some examples:
@@
These examples are used to help you clarify the expression form of instructional strategies, which are not entirely effective instructional strategies. You need to induce them yourself. Finally, please output in the following format: {“analysis”: xxx, “strategy”: xxx}, where ”analysis” is the reason why humans generate such instruction based on the dialogue history, and ”strategy” is just one induced instructional strategy. The expression should be concise, with no more than 20 words. Now please start thinking seriously. The results of your thinking are very important to me.
Table 6: The prompt for instructional strategy extraction.
Prompt for Instructional Strategy Abstraction
You are very good at generalizing a series of original instructional strategies into a high-level instructional strategy. High-level instructional strategies are general principles for thoughtfully, purposefully, and strategically generating instructions to obtain the desired information, create dialogues, stimulate thinking, or encourage others to share. To better reuse instructional strategies in new dialogue scenarios, you need to generalize a series of similar original instructional strategies into a higher-level instructional strategy. Each original instructional strategy is a specification of a high-level instructional strategy. We provide a list of similar original instructional strategies:{The List of Instructional Strategies}, and ask you to generalize them into a high-level instructional strategy. The high-level instructional strategy need to be a simple phrase. Next, please directly output the high-level instructional strategy: xxx.
Table 7: The prompt for instructional strategy abstraction.
Prompt for the User Simulator
You are an expert at generating instructions in human-AI conversations. The following is a dialogue between a user and an AI assistant. User statements start with [USER] and AI assistant statements start with [ASSISTANT]. To make it easier for you to ask for an insightful instruction, I will provide you a candidate of instructional strategies. Instructional strategies are general principles for thoughtfully, purposefully, and strategically generating instructions to obtain desired information, create conversation, stimulate thinking, or encourage others to share. You first need to choose an instructional strategy that fits the current flow of the dialogue from the candidate, and then generate a specific instruction based on the high-level instructional strategy and the dialogue history.  The candidate of instructional strategies is: {Instructional Strategies Candidate}. The dialogue history is: {Dialogue History}.
Table 8: The prompt for the user simulator.
@@
(4) The raised instruction may be a new instruction about the information related to the concept or entity contained in the answers from the ASSISTANT. The dialogue history is {Dialogue History}, the existing instructions are {Instruction List}, and the currently raised instruction is {Instruction}. Next, please carefully analyze whether the raised instruction is reasonable. Please be sure to output according to the following JSON format: {"analysis": "xxx", "result": "xxx"}, where “analysis” is the reason for whether it is reasonable, and “result” is "yes" or "no".
@@
Prompt for Data Evaluation
@@
The current instruction is: {Instruction}. Each evaluation metric counts for 10 points and you will rate the current instruction based on this dialogue history, with a high score representing better performance. You should carefully analyze the quality, and then give each score. Please be sure to output in the following JSON format: {"analysis": "[the quality in five evaluation dimensions]", "score": {"Appropriateness": "[an integer number between 1 and 10]", "Coherence": "[an integer number between 1 and 10]", "Depth": "[an integer number between 1 and 10]", "Insight": "[an integer number between 1 and 10]", "Diversity": "[an integer number between 1 and 10]"}}.
@@
A Constructed Multi-Turn Instructional Dialogue by ParrotChat

### v2

@@ 6 Related Work
Instructional Dialogue Construction.
Aligning LLMs with human expectations requires massive high-quality instructions. With the success of top-tier LLMs, it is feasible to use strong LLMs to automatically construct a variety of instructions, e.g., single-turn instructions Wang et al. (2023b); Wu et al. (2023); Anand et al. (2023); Honovich et al. (2023); Yu et al. (2023); Xu et al. (2023a) and multi-turn instructions Ding et al. (2023); Xu et al. (2023b); Li et al. (2023); Ji et al. (2023). The latest framework for constructing multi-turn instructions is leveraging two LLMs to interact: one simulating a user to pose instructions, and the other acting as a system agent to respond. Our proposed method is used to improve the performance of this framework of multi-turn instructional dialogue construction.
Instructions Generation.
The user simulator currently uses two implementations to generate instructions. One is to employ role-played LLMs to simulate humans through instruction prompts Xu et al. (2023b); Ding et al. (2023). The other is to fine-tune LLMs to learn instruction generation from real instructional dialogues Kong et al. (2023a); Sun et al. (2023). Besides, some works in similar tasks study how to ask high-quality questions, including question generation in open-domain dialogue systems Wang et al. (2018); Shen et al. (2021) and conversational question generation Gu et al. (2021); Li et al. (2022); Zeng et al. (2023). These works capture the underlying rules that guide the dialogue flows in different directions. In contrast, IDEAS explicitly models these underlying rules by introducing instruction strategies, to guide the dialogue flow in different directions.
@@
Some researches Ou et al. (2022); Bao et al. (2023) utilize a sequence of knowledge pieces (e.g., keywords, Wikipedia pages) to simulate the trajectory of dialogue flows for guiding open-domain dialogue generation. Compared to existing works, we are the first to leverage abstract rules, i.e., high-level instruction strategies, for dialogue flow guidance. We posit that LLMs have the capability to identify keywords and related information within extensive corpora. However, LLMs encounter difficulty in deriving underlying general rules. Thus, introducing high-level instruction strategies can potentially serve as a more effective direction, supporting the model in generating superior multi-turn dialogues.
@@
Our work is based on inductive learning Michalski (1983), which has shown promising results in various scenarios, including reasoning tasks Wang et al. (2023a); Zhu et al. (2023); Sun et al. (2024) and autonomous agents Park et al. (2023); Zhao et al. (2023). However, our work focuses on inducing high-level instruction strategies from human-machine dialogues and then applying them to guide instruction generation in new dialogue scenarios, which has not been explored in instructional dialogue construction.
@@ 7 Conclusion
This paper presents an inductive-deductive strategy reuse method, IDEAS, to generate diverse, in-depth instructions for building multi-turn instructional dialogues. Specifically, IDEAS first induces high-level instruction strategies from real dialogues via inductive reasoning. Afterward, IDEAS deductively applies these strategies to new dialogue scenarios, where the strategies serve as underlying rules to guide the dialogue flow in different directions, i.e., posing various instructions. Experimental results show that IDEAS can generate diverse and in-depth instructions. The constructed instructional dialogues can be used to improve the performance of the downstream chat model. We hope IDEAS will inspire more human-inspired methods to build higher-quality instructional dialogues.
@@
Due to limited computational and financial resources, we only select the fine-tuned LLM as the user simulator due to its lightweight and free. In fact, our designed IDEAS is also applicable to user simulators implemented by role-playing LLMs. Some researches Wang et al. (2023a); Zhu et al. (2023); Sun et al. (2024) propose similar insights, which have confirmed that using abstract rules can improve the performance of LLMs based on instruction prompts. In addition, we select the 13B model for fine-tuning. Recent research suggests that LLMs exhibit emerging capabilities when they expand beyond a certain threshold Wei et al. (2022a). Using IDEAS for user simulators based on larger LLM implementation may bring different effects. However, we do not have enough resources to conduct these relevant experiments. We welcome further researchers to study the benefits of abstract rules, i.e., instruction strategies, to user simulators of different scales and implementations. Besides, since this is the first study to propose an inductive-deductive strategy reuse method to propose diverse and in-depth instructions, our method design is relatively simple. From the experiment, the current IDEAS is also feasible and effective. We will optimize the design of IDEAS in the future.
Application Limitations.
When using IDEAS for building instructional dialogues with a specific domain, it is necessary to ensure that there are a certain number of human-machine dialogues in the relevant domain to extract the domain-dependent general strategies. This is also consistent with people’s cognition that it still takes a certain amount of experience to abstract general guiding rules in a specific domain. When adapting high-level strategies to specific domains, our designed ranker can select the set of appropriate strategies. These strategies may be domain-independent general strategy or domain-dependent general strategy. The domain-independent general strategy means that the instruction can be adapted in most domains, e.g.,“eliciting targeted information by asking context-specific questions to deepen understanding, expand knowledge, or facilitate decision-making”. The domain-dependent general strategy means that the strategy may only adapted to certain specific domains. For example, the strategy “emotional elicitation and empathetic inquiry” is more suitable for the domains “role-playing” and “chitchat”.
This paper chooses to extract strategies from ShareGPT data and selects opening lines from the domains covered by ShareGPT for testing. The top 10 most frequent strategies extracted from ShareGPT are shown in Table 8. We further conduct the following experiment, i.e., manually select 8 common domains in instructional dialogues to analyze the domain distribution of instructions in ShareGPT data and our constructed dialogue datasets. The distributions are shown in Table 6 and 7. We observe that ShareGPT contains almost no dialogues from other domains. All opening lines we use come from these 8 common domains. Besides, the domain distribution of our generated instructions in the ShareGPT scenario is very similar to that of the ShareGPT dataset. Further, we find that our generated instructions in the UltraChat scenario mainly cover two domains: theoretical knowledge and applied knowledge. This is mainly because UltraChat contains plenty of knowledge-based instructional dialogues. We also find that ShareGPT also includes the chitchat dialogues, not just task-completing instructions. Chitchat can actually be regarded as a special kind of instruction since chitchat-dialogue generation can also introduce instruction strategies to guide.
@@
We implement instruction strategy extraction and generalization, along with quality control by prompting GPT-4. We also choose GPT-4 as the system agent. However, GPT-4 is only accessed through an interface. The mechanisms behind these interfaces may change at any time, so the constructed dialogues from different periods may change.
@@
Given that GPT-4 is trained on online data, it is conceivable that GPT-4 may encode pervasive biases that perpetuate stereotypes, discrimination, or marginalization of specific languages or communities. This results in potentially toxic and harmful answers from GPT-4 as the system agent. In addition, our evaluation process involves manual intervention by two professional annotators. Each annotator is compensated \$$\$0.2$ per instance.
Appendix A IDEAS
Coding
Applied
Knowledge
Role
Playing
Theoretical
Knowledge
Brainstorm Reasoning Chitchat Math Others
ShareGPT 34.10% 27.99% 14.01% 7.13% 6.73% 5.23% 3.95% 0.77% 0.09%
Table 6: The domain distribution of ShareGPT data that are used to extract instruction strategies.
Coding
Applied
Knowledge
Role
Playing
Theoretical
Knowledge
Brainstorm Reasoning Chitchat Math Others
ShareGPT 27.84% 23.06% 18.51% 14.85% 3.70% 4.09% 4.57% 1.31% 0.00%
UltraChat 0.73% 38.72% 2.96% 50.28% 1.13% 5.04% 1.10% 0.03% 0.00%
Table 7: The domain distribution of opening lines in ShareGPT and UltraChat dialogue scenarios.
A.1 Model Training
@@
The BERT-based ranker may have limited performance. Therefore, this may raise a question: why not train a stronger ranker to select the top-$1$ strategy or randomly select a strategy from the top-$k$ strategies, i.e., the candidate, to avoid further selection by the user simulator? This is mainly due to the following reasons: (1) There are generally multiple appropriate instruction strategies for a given dialogue history, so it is unreasonable to choose the top-$1$ strategy. (2) It is reasonable for the user simulator to sample one from multiple appropriate strategies in the candidate. (3) Limited performance means that not every strategy in the candidate is appropriate, so random selection from the candidate is also unreasonable. (4) The strategy selection of the user simulator is better than that of the ranker, whose secondary selection can select more appropriate strategies. (5) It is easy for the user simulator to select a strategy before generating instructions, but training a stronger ranker will greatly increase the complexity of IDEAS. Overall, this lightweight model is sufficient for a coarse selection of candidates.
Appendix B Experimental Details
B.1 Data
The basic data used for all experiments are collected by Sun et al. (2023). For instruction generation, the human-machine instructional dialogues comes from ShareGPT  [^4] https://huggingface.co/datasets/shareAI/ShareGPT-Chinese-English-90k  [^5] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna
@@ B.1 Data
to obtain instruction strategies, along with training the user simulator and the ranker in the induction stage. The dialogues are further cleaned by using the provided method  [^6] https://huggingface.co/datasets/shareAI/ShareGPT-Chinese-English-90k/blob/main/shareGPT/filter_data.py. Finally, we obtain $56,929$ dialogues. Subsequently, we process each multi-turn dialogue $\{\boldsymbol{q}_{0},\boldsymbol{a}_{0},\dots,\boldsymbol{q}_{T},\boldsymbol{a}_{T}\}$ into $T$ pairs of dialogue histories and instructions, e.g., <$<\{\boldsymbol{q}_{0},\boldsymbol{a}_{0},\dots,\boldsymbol{a}_{t-1}\},\boldsymbol{q}_{t}>$. We further obtain $211,495$ pairs and the corresponding original instruction strategies. Afterward, we achieve $1,593$ high-level instruction strategies via instruction strategy abstraction. We count the top 10 most frequent instruction strategies extracted from ShareGPT data and give the number and ratio of occurrences in Table 8. Furthermore, we randomly sample $200$ pairs of dialogue histories and instructions and employ three human annotators to evaluate the appropriateness of original and high-level instruction strategies. About $95.67\%$ original instruction strategies are accepted by at least two annotators, and about $95\%$ high-level instruction strategies are accepted. Similarly, we evaluate the accuracy of the original strategy abstraction, and at least two annotators accept about $94.33\%$ corresponding high-level strategies.
In the deduction stage, we randomly sample $10,000$ dialogues from unused ShareGPT and UltraChat datasets Sun et al. (2023) respectively, and then intercept the first-turn dialogues $\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}\textquoteright_{0}\}$ as opening lines. We set the maximum interaction to $10$ rounds and the maximum of regeneration in quality control to $5$. The average number of regeneration is $0.69$. We also evaluate the proportion of instruction strategies selected by the user simulator from the candidates, which is $97.76\%$. Besides, we evaluate the appropriateness of the selected strategies by the user simulator, and the annotators accept about $93.82\%$ selected instruction strategies. Furthermore, we manually evaluate the correctness of quality control, and at least two annotators accepted $93.33\%$ of the judgment results. After achieving our constructed instructional dialogues $\mathcal{D}_{chat}$, we clean dialogues from $\mathcal{D}_{chat}$ by removing subsequent rounds starting from the round in which the error content appears. The error content contains the empty instructions and the incorrect answers caused by exceptions in calling the interface API of the system agent.
B.2 Implementation Details
@@
To validate the effectiveness of IDEAS, it is necessary to ensure the quality of each module when implementing IDEAS. To this end, we choose the top-tier model currently, i.e., GPT-4. Therefore, we prompt gpt-4-1106-preview for instruction strategy extraction and abstraction, along with quality control. We set the temperature to 00, the presence_penalty to $0.6$, and the frequency_penalty to 00. For strategy clustering, the threshold $\epsilon$ is set to $0.5$. The entire cluster process is implemented by Faiss vectorstore Johnson et al. (2019). For ranking, we set the threshold $\eta$ to $0.5$. We limit the size of the candidates to 50 due to the context length limitation. When generating instructions, we use FasterTransformer  [^7] https://github.com/NVIDIA/FasterTransformer for inference acceleration. And the temperature is set to $0.7$, the top-p is $0.9$, and the max new tokens is $96$. The average inference time for one instruction is approximately $0.2$ second. Besides, we follow existing work and employ the top-tier model, i.e. gpt-4-1106-preview as the system agent, where the parameters are default.
@@
Rank High-level Instruction Strategy Percentage Count
1 Eliciting targeted information by asking context-specific questions to deepen understanding, expand knowledge, or facilitate decision-making $3.95\%$ 8337
2 Requesting specific, practical examples and real-world applications to clarify and deepen understanding of theoretical concepts, ensure comprehensive comprehension, and facilitate the application of knowledge in various contexts $3.92\%$ 8273
3 Requesting elaboration, clarification, and additional information to enhance understanding and obtain comprehensive context $3.32\%$ 7007
4 Requesting detailed, step-by-step information or guidance tailored to specific tasks, scenarios, or contexts to ensure accurate understanding and implementation $2.72\%$ 5740
5 Seeking practical and actionable solutions through clarification, troubleshooting, and expert guidance for identified problems within various contexts $2.57\%$ 5424
6 Clarifying Concepts and Terminology $2.22\%$ 4685
7 Encourage detailed elaboration and comprehensive understanding through open-ended, context-specific inquiry $1.92\%$ 4052
8 Strategic Inquiry for Expert Insight on Complex Topics $1.42\%$ 2997
9 Proposing and soliciting feedback for enhancements, alternatives, and improvements to existing solutions, methods, or ideas $1.42\%$ 2997
10 Applying and adapting knowledge and concepts to diverse contexts and scenarios for enhanced understanding and problem-solving $1.37\%$ 2891
Table 8: The top 10 most frequent instruction strategies in ShareGPT data.
Prompt for Instruction Strategy Extraction
You are very good at learning from real-human dialogues about how they generate high-quality instructions. The dialogue history is: {Dialogue History}. The next instruction generated by the human is: {Instruction}. Please analyze why humans generate such instruction based on the above dialogue history, and give the instruction strategy you have learned. Good instruction strategies need to meet the following requirements:
(1) Instruction strategies should be expressed in simple phrases;
(2) Instruction strategies should not include specific information from the dialogue history. To better assist you in inducing instruction strategies, I will provide you with some examples:
@@
These examples are used to help you clarify the expression form of instruction strategies, which are not entirely effective instruction strategies. You need to induce them yourself. Finally, please output in the following format: {“analysis”: xxx, “strategy”: xxx}, where ”analysis” is the reason why humans generate such instruction based on the dialogue history, and ”strategy” is just one induced instruction strategy. The expression should be concise, with no more than 20 words. Now please start thinking seriously. The results of your thinking are very important to me.
Table 9: The prompt for instruction strategy extraction.
Prompt for Instruction Strategy Abstraction
You are very good at generalizing a series of original instruction strategies into a high-level instruction strategy. High-level instruction strategies are general principles for thoughtfully, purposefully, and strategically generating instructions to obtain the desired information, create dialogues, stimulate thinking, or encourage others to share. To better reuse instruction strategies in new dialogue scenarios, you need to generalize a series of similar original instruction strategies into a higher-level instruction strategy. Each original instruction strategy is a specification of a high-level instruction strategy. We provide a list of similar original instruction strategies:{The List of Instruction Strategies}, and ask you to generalize them into a high-level instruction strategy. The high-level instruction strategy need to be a simple phrase. Next, please directly output the high-level instruction strategy: xxx.
Table 10: The prompt for instruction strategy abstraction.
Prompt for User Simulator
You are an expert at generating instructions in human-AI conversations. The following is a dialogue between a user and an AI assistant. User statements start with [USER] and AI assistant statements start with [ASSISTANT]. To make it easier for you to ask for an insightful instruction, I will provide you a candidate of instruction strategies. Instruction strategies are general principles for thoughtfully, purposefully, and strategically generating instructions to obtain desired information, create conversation, stimulate thinking, or encourage others to share. You first need to choose an instruction strategy that fits the current flow of the dialogue from the candidate, and then generate a specific instruction based on the high-level instruction strategy and the dialogue history.  The candidate of instruction strategies is: {Instruction Strategies Candidate}. The dialogue history is: {Dialogue History}.
Table 11: The prompt for user simulator.
@@
(4) The raised instruction may be a new instruction about the information related to the concept or entity contained in the answers from the ASSISTANT. The dialogue history is {Dialogue History}, the existing instructions are {Instruction List}, and the currently raised instruction is {Instruction}. Next, please carefully analyze whether the raised instruction is reasonable. Please be sure to output according to the following JSON format: “analysis”: “xxx”, “result”: “xxx”}, where “analysis” is the reason for whether it is reasonable, and “result” is “yes” or “no”.
@@
Prompt for Instruction Evaluation
@@
The current instruction is: {Instruction}. Each evaluation metric counts for 10 points and you will rate the current instruction based on this dialogue history, with a high score representing better performance. You should carefully analyze the quality, and then give each score. Please be sure to output in the following JSON format: {“analysis”: “[the quality in five evaluation dimensions]”, “score”: {“Appropriateness”: “[an integer number between 1 and 10]”, “Coherence”: “[an integer number between 1 and 10]”, “Depth”: “[an integer number between 1 and 10]”, “Insight”: “[an integer number between 1 and 10]”, “Diversity”: “[an integer number between 1 and 10]”}}.
@@
A Constructed Multi-Turn Instructional Dialogue by Parrot-Ask